# Bench4Merge: A Comprehensive Benchmark for Merging in Realistic Dense Traffic with Micro-Interactive Vehicles

### v1

@@ IV-B Effectiveness of Micro-Controllable Vehicles Model
We then demonstrate that the vehicles are capable of reflecting the micro-level characteristics. To this end, we test the environment 100 times, during which we analyzed the differences in key metrics among three vehicle styles we configured. As shown in Table. III, The average distance between friendly and offensive vehicles in the environment is 5.31m and 3.45m, respectively, which closely aligns with 5.35m and 4.07m from the dataset. This demonstrates that our model effectively captures the micro behaviors.
@@
Avg.dhw (m) 4.07 3.45 5.35 5.31 6.49 6.69
Avg.speed (m/s) 2.96 2.31 2.95 2.21 2.03 1.51
Avg.acc_x (m/s) 0.17 -0.37 0.19 -0.31 0.15 -0.53
Avg.acc_x_std 0.31 1.36 0.41 0.81 0.09 0.65
Avg.acc_y (m/s) -0.07 -0.21 0.08 0.05 0.04 -0.08
Avg.acc_y_std 0.12 0.08 0.05 0.07 0.04 0.06
Avg.offset (m) -0.29 -0.26 -0.13 -0.14 0.04 0.04
We implemented a RL-based method [38] within our environment, the average scores and success rates are lower compared to those in the previous environment as shown in Table. IV. Indicating that previously rule-defined main road vehicles lack the micro-level behaviors of real-world traffic participants, resulting in overly simplistic scenarios. Notably, the vehicles exhibited varying success rates across the three different initial scenarios in our environment, underscoring the significance of our initial environment classification (Section III-B).
@@
IDM-based Environment 85% 6.16
Bench4Merge Highly Dense 55% 5.41
Medium Dense 35% 3.26
Lower Dense 50% 5.05
@@ IV-C Evaluation of Representative Methods
As shown in Figure 7, we can draw the following conclusions: First, we observed that none of the methods performed well across all aspects. The RL-based method achieved the highest score but exhibited issues with sharply acceleration changes [38]. In contrast, the Imitation Learning method had the lowest score but demonstrated advantages in passing speed [39]. Second, the suggestion “Enhance awareness” was the most frequently given for optimization-based methods [40, 41], indicating a significant lack of interaction with surrounding vehicles. On the other hand, “Smooth Acceleration” was the most common suggestion for neural network-based methods [3, 38, 39], highlighting a deficiency in comfort. Finally, Bench4Merge also revealed issues that had not been identified in previous work, Hierarchical RL have shown significant advantages over Spatio-temporal Planner [41] in previous work [3]. However, in Bench4Merge, the average score reflects only a marginal advantage and is lower than that of RL method. This is because prior metrics did not account for comfort, whereas Bench4Merge offers a more comprehensive evaluation.
@@ V Conclusion
In this work, we introduced the Bench4Merge for dense merging scenarios, aimed at evaluating the performance of motion planning methods comprehensively. We extracted initial scenarios from large-scale data and classify them accordingly and trained surrounding vehicles to exhibit realistic and diverse interaction behaviors. Finally, we redefine the evaluation mechanism by leveraging an LLM to score and analyze each sample. This evaluation approach overcoming the limitations of previous evaluation methods. From the analysis, we have distilled valuable analysises from various methods, and have open-sourced all environments and algorithms. Future work will focus on generating 3D environments for dense interaction scenarios with scene rendering, catering to the closed-loop testing needs of the increasingly prominent multimodal end-to-end autonomous driving systems.

### v3

@@ IV-B Effectiveness of Micro-Controllable Vehicles Model
Now we demonstrate that the vehicles are capable of reflecting the micro-level characteristics. To this end, we test the environment 100 times, during which we analyzed the differences in key metrics among three vehicle styles we configured. As shown in Table. III, The average distance between friendly and offensive vehicles in the environment is 5.81m and 3.82m, respectively, which closely aligns with 5.35m and 4.07m from the dataset. This demonstrates that our model effectively captures the micro behaviors.
@@
Avg.dhw (m) 04.07 03.82 05.35 05.81 06.49 07.73
Avg.speed (m/s) 02.96 02.40 02.95 02.15 02.03 01.62
Avg.acc_x (m/s) 00.17 -0.11 00.19 -0.08 00.15 -0.26
Avg.acc_x_std 00.31 01.91 00.41 01.56 00.09 01.43
Avg.acc_y (m/s) -0.07 -0.07 00.08 00.03 00.04 00.01
Avg.acc_y_std 00.12 00.09 00.05 00.06 00.04 00.06
Avg.offset (m) -0.29 -0.23 -0.13 -0.14 00.04 00.02
We implemented a RL-based method [40] within our environment, the average scores and success rates are lower compared to those in the previous environment as shown in Table. IV. Indicating that previously rule-defined main road vehicles lack the micro-level behaviors of real-world traffic participants, resulting in overly simplistic scenarios. Notably, the vehicles exhibited varying success rates across the three different initial scenarios in our environment, underscoring the significance of our initial classification (Section III-B).
@@
IDM-based Environment 85% 5.96
Bench4Merge Highly Dense 55% 4.99
Medium Dense 35% 3.48
Lower Dense 50% 4.42
@@ IV-C Evaluation of Representative Methods
As shown in Figure 7, we can draw the following conclusions: First, we observed that none of the methods performed well across all aspects. The RL-based method[40] achieved the highest score but exhibited issues with sharp acceleration changes. In contrast, the Imitation Learning method [41] had the lowest score but demonstrated advantages in passing speed. Second, the suggestion “Enhance awareness” was the most frequently given for optimization-based methods [42, 43], indicating a significant lack of interaction with surrounding vehicles. On the other hand, “Smooth Acceleration” was the most common suggestion for neural network-based methods [3, 40, 41], highlighting a deficiency in comfort. Finally, Bench4Merge also revealed issues that had not been identified in previous work, Hierarchical RL have shown significant advantages over Spatio-temporal Planner [43] in previous work [3]. However, in Bench4Merge, the average score reflects only a marginal advantage and is lower than that of RL method. This is because prior metrics did not account for comfort, whereas Bench4Merge offers a more comprehensive evaluation.
@@ V Conclusion
In this work, we introduced the Bench4Merge for dense merging scenarios, aimed at evaluating the performance of motion planning methods comprehensively. We extracted initial scenarios from large-scale data and classify them accordingly and trained surrounding vehicles to exhibit realistic and diverse interaction behaviors. Finally, we redefine the evaluation mechanism by leveraging an LLM to score and analyze each sample. This evaluation approach overcoming the limitations of previous evaluation methods. From the analysis, we have distilled valuable analyses from various methods, and have open-sourced all environments and algorithms. Future work will focus on generating 3D environments for dense interaction scenarios with scene rendering, catering to the closed-loop testing needs of the increasingly prominent multimodal end-to-end autonomous driving systems.
ACKNOWLEDGEMENTS
This work is supported by Wuxi Research Institute of Applied Technologies, Tsinghua University under Grant 20242001120 and Didi Chuxing.
Appendix A: Training Data Construction
Figure 8: The evaluation of real-world autonomous driving platform based on the LLM provides valuable guidance for the planning
Based on the definitions used in existing work for dense traffic scenes, we set the following requirements to filter the data: the Target Vehicle and the Leading Vehicle must be on the same lane, with a distance between them no greater than 10 meters, and their speeds must range from 1 $m/s$ to 5 $m/s$. Additionally, there must be interacting vehicles within a 5-meter range both in front and behind the Target Vehicle in the adjacent lanes.
Finally, our data composition is as follows: The DJI dataset contains 270,000 samples, while the nuPlan and ExiD datasets each contain 100,000 samples, we also used 100,000 samples collected from the DiDi test vehicle. Each sample is composed of four components: the Target vehicle, the Leading Vehicle, the Interactive Vehicle, and the map information, as shown in the Fig. 5. The data for Target Vehicle is represented as an 11-dimensional vector, The dataset includes the following vehicle parameters: lateral position, longitudinal position, heading angle, lateral speed, longitudinal speed, lateral acceleration, longitudinal acceleration, total acceleration, steering angle, distance to the preceding vehicle (Thw), offset distance from the lane centerline (Offset), label (as introduced in SectionIII-C), and length.
Appendix B: Example of LLM-Based Evaluation
We have deployed the evaluation mechanism onto our test vehicle. Fig. 8 illustrates one evaluation process of the system. The input to the LLM consists of four main components: First, all the data related to the ego vehicle’s merging process, including average speed, total time, the coordinates of the merging point, acceleration fluctuations, and the average and minimum distances maintained with adjacent vehicles. Second, the main lane traffic data, which includes the average speed of vehicles on the main road. Third, prior knowledge set based on experience, which includes the definition of comfortable acceleration and the range of efficient passing times. Fourth, the status of the ego vehicle, which can be selected from either relax or hurry.